# M. C. Rodriguez

### v5

@@ Abstract
We present a detailed study of the usual scalar potential of the Minimal Supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ Model. In this article, we will consider the case in wich all the usual neutral scalars fields of this model obtain vacuum expectation values. When we allow in our superpotential, only interactions that respect the invariance of the quantum number, ${\cal F}$. We obtain, as in the minimal $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ model without supersymmetry, a Majoron. However, by allowing the violation of this quantum number, we can remove the Majoron in a mechanism similar to that recently presented in the Supersymmetric version of the scheme of Gelmini-Roncadelli. The mass spectrum obtained is in agreement with current experimental data.
@@ 1 Introduction
In the context of the Standard Model (SM) [1, 2], we introduce only left-handed neutrinos [^1] We wish to highlight a SM hypothesis, not to introduce right-handed neutrinos, $\left[\nu_{iR}\sim\left({\bf 1},{\bf 1},0\right)\right]$. Those right-handed neutrinos are known as sterile neutrinos [3, 4]., in the following representations of $SU(2)_{L}$ [5, 6]
$L_{i}$ $=$ $\left(\begin{array}[]{c}\nu_{iL}\\ l_{iL}\end{array}\right)\sim({\bf 1},{\bf 2},-1),\,\ l_{iR}\sim({\bf 1},{\bf 1},-2).\,$ (3)
We show in parenthesis the transformations properties under the respective factors $(SU(3)_{C},SU(2)_{L},U(1)_{Y})$, and $i=1,2,3,$ is family index. We are omitting the quark sector, and they can be found at [7, 8, 9].
In the SM, we have to introduce the following scalar field $\phi$ in the doublet representation of the $SU(2)_{L}$
$\phi=\left(\begin{array}[]{c}\phi^{+}\\ \phi^{0}\end{array}\right)\sim({\bf 1},{\bf 2},+1),\,\ \Rightarrow\,\ \tilde{\phi}\equiv\left[\left(i\sigma_{2}\right)\phi^{*}\right]=\left(\begin{array}[]{c}\left(\phi^{0}\right)^{*}\\ -\phi^{-}\end{array}\right)\sim({\bf 1},{\bf{\bar{2}}},-1).$ (8)
This scalar fields get the following vaccum expectation value (VEV)
@@
The charged leptons obtain their masses, via the following Yukawa coupling [^2] As well as for all other particles of the model, for more details see [7, 8, 9].
${\cal L}^{Y}_{MP}$ $=$ $g^{l}_{ij}\left[\left(\bar{L}_{iL}\phi\right)l_{jR}+hc\right].$ (14)
The left-handed neutrinos fields ($\nu_{iL}$) are massless at all perturbative order.
The scalar potential of the SM is defined in the following way [10]
@@
Both the $\mu$-parameter and the $\lambda$-parameter are arbitrary, that is, they are free parameters of the SM. For $\mu$-parameter, we can have the following two possibilities [10]
  • $\mu^{2}>0:$ The gauge symmetry $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ is not break;
  • $\mu^{2}<0:$ We break the $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ gauge symmetry to $SU(3)_{C}\times U(1)_{em}$,
and the minimum of the potential, when $\mu^{2}<0$, is written in the following way
$v$$\equiv$ $\sqrt{\frac{-\mu^{2}}{\lambda}}=\frac{1}{\sqrt{\sqrt{2}G_{F}}}=246\,\ {\mbox{G}eV},$
$G_{F}$ $=$ $\left(1.16639\pm 0.00002\right)\times 10^{-5}\,\ {\mbox{G}eV}^{-2}\approx\frac{1}{(293\,\ GeV)^{2}}.$ (16)
@@
$V(\phi^{\prime 0})$ $=$ $-\frac{\lambda}{4}v^{4},\,\ |\phi^{\prime 0}|=\sqrt{\frac{-\mu^{2}}{2\lambda}}\equiv\frac{v}{\sqrt{2}}>0.$ (18)
The field $(h^{0})$ [^3] The real part of $\phi^{0}$, see Eq.(17). is CP even field and get the following mass  [^4] A nice review about this interesting subject can be found in [11].
@@
We must emphasize, the SM do not predict any value for this mass, in fact it can get any values from 00 until a superior limit given by $M_{h^{0}}<1$ TeV [12, 13]. By another hand, the field $(G^{0})$ [^5] The imaginary part of $\phi^{0}$, see Eq.(17), and it is known as Goldstone boson [14]. is CP odd field and it is massless field. This field $(G^{0})$, will subsequently generate mass to the massive gauge boson $Z^{0}$, this mechanism is known as Brout-Englert-Higgs mechanism [11, 15].
The mass of the light Higgs boson $h^{0}$ was discovered experimentally by the ATLAS (A Toroidal LHC Apparatus) and CMS (Compact Muon Solenoid) collaborations at Large Hadron Collider (LHC) [16, 17]
@@
After breaking the gauge symmetry, we obtain the following two massive gauge bosons [7, 8]
@@
$M^{2}_{Z^{0}}$ $=$ $\left(\frac{g^{2}+(g^{\prime})^{2}}{4}\right)v^{2}=\frac{g^{2}v^{2}}{4\cos^{2}\theta_{W}}=\frac{M^{2}_{W^{\pm}}}{\cos^{2}\theta_{W}},$
where $\theta_{W}$ is the Weinberg angle and the numerical was taken from [18]. The Weinberg angle is defined as follows  [^6] Where, $g$ and $g^{\prime}$ are the gauge constant of the group $SU(2)_{L}$ and $U(1)_{Y}$, respectively.
@@
The average values for the gauge bosons masses are [16, 18]
@@
Today we know neutrinos are massive particles [19, 20, 21, 22]. A priori, to solve this problem, we can introduce the following Majorana mass term for the left-handed neutrinos [^7] If we introduce sterile neutrinos in SM, $\nu_{iR}$, we can write the following Dirac mass term for neutrinos $y^{\nu}_{ij}\left[\left(\bar{L}_{iL}\tilde{\phi}\right)\nu_{jR}+hc\right]$, where $\tilde{\phi}$ is defined in our Eq.(17). This mass term conserve lepton number defined in our Eq.(26). [21]
@@
where $B$ is the baryon number and $L_{e},L_{\mu}$ and $L_{\tau}$ are the lepton number of each lepton flavor  [^8] It means $L_{e}$ is lepton number of electrons and in similar way for $L_{\mu}$, $L_{\tau}$ is the lepton number of muons and of taus, respectively..
An interesting way to generate the Eq.(25), is to add the following term for the lagrangian of SM [9]
$\overline{L^{c}_{iL}}L_{jL}\sim\left({\bf 1},{\bf 1\oplus 3},-2\right);\,\ L^{c}_{iL}=\left[\left(i\sigma_{2}\right)L^{*}_{iL}\right]=\left(\begin{array}[]{c}l^{c}_{iL}\\ -\nu^{c}_{iL}\end{array}\right)\sim\left({\bf 1},{\bf{\bar{2}}},+1\right).$ (27)
Therefore, to provide mass to neutrinos, we need to introduce new fields within the SM context. We can add to the SM singlets and also the triplets, of scalars or fermions, in representation of $SU(2)_{L}$. When we introduce
@@
We can break the lepton number symmetry [^9] The total lepton number $L$ is defined in terms of $L_{e},L_{\mu}$ and $L_{\tau}$, see Eq.(26), in the following way $L=L_{e}+L_{\mu}+L_{\tau}$. when we introduce the new scalar $\Delta$ in the triplet representation, see Eq.(27), of the group $SU(2)_{L}$
$\Delta=\left(\begin{array}[]{cc}\Delta^{0}&\frac{h^{+}}{\sqrt{2}}\\ \frac{h^{+}}{\sqrt{2}}&H^{++}\end{array}\right)\sim\left({\bf 1},{\bf 3},+2\right),\,\ \langle\Delta\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}V_{\Delta}&0\\ 0&0\end{array}\right).$ (28)
@@
${\cal L}_{\Delta,L}$ $=$ $g^{\nu}_{ij}\left[\left(\overline{L^{c}_{iL}}\Delta L_{jL}\right)+hc\right]$
$=$ $g^{\nu}_{ij}\left[-\overline{\nu^{c}_{iL}}\nu_{jL}\Delta^{0}+\left(\overline{l^{c}_{iL}}\nu_{jL}-\overline{\nu^{c}_{iL}}l_{jL}\right)\frac{h^{+}}{\sqrt{2}}+\overline{l^{c}_{iL}}l_{jL}H^{++}+hc\right].$
Those scalars $\Delta^{0},\,\ h^{+}$ and $H^{++}$ are known as bileptons [38]. The first term in the above equation is exactly our Eq.(25), where we can identify
@@
It is know as Gelmini-Roncadelli scheme [37]. The double beta decay experiments, is defined in the following way
@@
can show if the neutrinos are Majorana particles [39] or Dirac particles [40], more information on this subject can be seen at [5, 6, 21, 41, 42, 43].
@@
Using the new scalar triplet, defined in our Eq.(28), we can also explain the new measurement for the $W$-boson mass [44]
$\left(M_{W^{\pm}}\right)_{\rm CDF}=\left(80.4335\pm 0.0094\right)\,\ GeV,$ (34)
because the mases of the gauge bosons in the scheme of Gelmini-Roncadelli are written as follows [45, 46, 47]
$M^{2}_{W^{\pm}}$ $=$ $\frac{g^{2}}{4}\left(v^{2}+2V^{2}_{\Delta}\right)\frac{g^{2}v^{2}}{4}\left(1+2R\right),\,\ R=\frac{V^{2}_{\Delta}}{v^{2}},$
@@
We realize that we can explain the new measurement for the $W$-boson mass if $V_{\Delta}\approx 10$ GeV, as can be easily seen in our Fig.(1).
Figure 1: We show the prediction about the behaviour for the gauge boson mass, $M_{W^{\pm}}\equiv M_{W}$, in the scheme of Gelmini-Roncadelli as a function of the Triplet VEV $V_{\Delta}$ (black curve) and we take $v=246$ GeV. The straight lines in red, green and blue are the values shown in our Eqs.(LABEL:wsmval,23,34), respectively.
@@
Now if we make the shift defined in our Eq.(17), together with
@@
After carrying out the shift, defined in our Eqs.(17,38), we obtain the following mass matrices for the neutral scalars
@@
Where we get the following analytical results $det\left(M^{2}_{R}\right)\neq 0$ and $Tr\left(M^{2}_{R}\right)\neq 0$, therefore we have two massive scalars. The lightest Higgs, we will represent as ($h^{0}$), and we also have another heavy Higgs, which we will denote as ($H^{0}$). The mass of the lightest scalar, is the same as the Higgs boson of SM compare with Eq.(19), is [^10] The complete expression for the mass of these scalars can be found in [51]. The mass of heavier Higgs, $H^{0}$, we will present below.
@@
The mass $M_{h^{0}}$ as function of the $\lambda_{1}$-parameter is shown in our Fig.(3). Where, we can see that we can get the central data of the lightest Higgs mass if we assume that $\lambda_{1}\approx 0.1295$.
@@
We get two Goldstone bosons, because $det\left(M^{2}_{I}\right)=0$ and $Tr\left(M^{2}_{I}\right)=0$, one of them is the same $G^{0}$ obtained in SM and necessary for giving mass to $Z^{0}$. The extra Goldstone boson is the Majoron, $M^{0}\equiv I_{\Delta}$ [37]. The Majoron, $M^{0}$, can be detected experimentally through the following process [41, 42, 43]
@@
The existence of a non-massive Majoron, $M^{0}$, is excluded by experimental data [^11] Decay of the $Z^{0}$ boson. [52]. One possible solution for this problem, is to make the Majoron, $M^{0}$, invisible to the leptons and quarks, but we will not use this solution in this article.
In order to remove the non-massive Majoron, there are another three possible solutions [50]
@@
Where $V_{G}(\phi,\Delta)$ is defined in our Eq.(37) and $\tilde{\phi}$ is defined in our Eq.(17). $\langle\Delta\rangle$ give mass only for neutrinos, review our Eq.(30), then it has tiny VEV. In the scheme of Ma, our Eq.(52) become
@@
The mass spectrum, we get beyond the lightest Higgs, $h^{0}$, given in Eq.(48), we obtain the heav Higgs field, $H^{0}$, and its mass is given by
@@
Figure 4: We show the prediction about the behaviour for the Higgs, $M_{H^{0}}$, in the scheme of Ma, see our Eq.(53), as a function of the Triplet VEV $V_{\Delta}$, and $v=246$ GeV. The values of $\mu\equiv\mu_{M}$ shown in the box in several colors. $M^{exp}_{H^{0}}$ is the lower experimental values given at our Eq.(54) in orange line.
@@
therefore $M_{A^{0}}>M_{H^{0}}$. We show $M_{A^{0}}$ as function of, the VEV, $V_{\Delta}$ in our Fig.(5) and again our results for the mass of $M_{A^{0}}$ is in agreement experimental limits for the pseudoscalar, $A^{0}$, is [53]
@@
if $V_{\Delta}\leq 3$ GeV and $\mu\geq 10^{0}$ GeV. A numerical example, when $V_{\Delta}=0.1$ GeV and $\mu=0.1$ GeV we obtain
$M_{H^{0}}=654.150\,\ GeV,\,\ M_{A^{0}}=777.921\,\ GeV,$ (57)
they are shown by the green lines in our Figs.(4,5).
Figure 5: We show the prediction about the behaviour for the pseudoscalar, $M_{A^{0}}$, in the scheme of Ma, see our Eq.(55), as a function of the Triplet VEV $V_{\Delta}$, and $v=246$ GeV. The values of $\mu\equiv\mu_{M}$ shown in the box in several colors. $M^{exp}_{A^{0}}$ is the lower experimental values given at our Eq.(56) in orange line.
@@
We notice, the charged scalars are heavier in the case $\lambda_{5}<0$ than when $\lambda_{5}>0$. Our numerical results satisfy the following experimental constraints $M_{h^{\pm}}>155$ GeV [16] and $M_{h^{\pm\pm}}>445$ GeV [56]. As a numerical example when $V_{\Delta}=0.1$ GeV and $\mu=0.1$ GeV, we get the values for the charged scalars
  • $\lambda_{5}>0$; $\lambda_{5}=0.3$
    $M_{h^{\pm}}=650.672\,\ GeV,\,\ M_{h^{\pm\pm}}=647.175\,\ GeV.$ (61)
  • $\lambda_{5}<0$; $\lambda_{5}=-0.3$
    $M_{h^{\pm}}=657.610\,\ GeV,\,\ M_{h^{\pm\pm}}=661.052\,\ GeV.$ (62)
Those values were take from our Figs(6,7).
Figure 6: We show the prediction about the behaviour for the masses, $M_{H^{0}}$ (blue curve), and $v=246$ GeV. The values of $\mu\equiv\mu_{M}$ shown in the box in several colors. The $M_{h^{\pm}}$ (green curve) and $M_{h^{\pm\pm}}$ (red curve) when $\lambda_{5}>0$. The values contained in this graph are in agreement with the analytical results obtained by Eqs.(60,61).
Figure 7: We show the prediction about the behaviour for the masses $M_{h^{\pm}}$ (green curve) and $M_{h^{\pm\pm}}$ (red curve) when $\lambda_{5}<0$ and $M_{H^{0}}=654.150$ GeV. The values contained in this graph are in agreement with the analytical results obtained by Eqs.(59,62).
Supersymmetry, it is more known as SUSY, arose in theoretical papers more than 30 years ago independently by Golfand and Likhtman [57], Volkov and Akulov [58] and Wess and Zumino [59]. The supersymmetry algebra was introduced in [57]  [^13] They also constructed the first four-dimensional field theory with supersymmetry, (massive) quantum electrodynamics (SQED) of spinors and scalars. We will call it as Golfand-Likhtman model as suggest in [60].. The Volkov-Akulov article, however, deals only with fermions and its name is “Is the Neutrino a Goldstone Particle? [^14] It is the base of supergravity.. For interested people interested in the history of supersymmetry see [61, 62] and for the earliest Supersymmetric Standard Model see [63, 64], while anyone interested, how to work with supersymmetry see the great review article [65].

### v6

@@ Abstract
We present a preliminar study of the scalar potential of the usual scalars in the Minimal Supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ Model. We will consider the case in which all the usual neutral scalars fields of this model obtain vacuum expectation values. When we allow in our superpotential, only interactions that respect the invariance of the quantum number, ${\cal F}\equiv B+L$, where $B$ is the baryon number while $L$ is the total lepton number. We obtain, as in the minimal $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ model without supersymmetry, a Majoron. However, by allowing the violation of this quantum number, we can remove the Majoron in a mechanism similar to that recently presented in the Supersymmetric version of the scheme of Gelmini-Roncadelli. The mass spectrum obtained is in agreement with current experimental data.
@@ 1 Introduction
In the Standard Model (SM) [1, 2], the relationship among the isospin $I_{3}$, the hypercharge $Y$ and the eletromagnetic charge $Q$ is defined by the Gell-Mann-Nishijima equation, which is defined as
$Q=I_{3}+Y.$ (1)
In the SM, we introduce only left-handed neutrinos [^1] We do not to introduce right-handed neutrinos, $\left[\nu_{iR}\sim\left({\bf 1},{\bf 1},0\right)\right]$, in agreement with the two component theory of neutrino theory of Landau [3], Lee and Yang [4] and Salam [5] from 1957, later the helicity of neutrinos was measured experimentally in 1958 and it confirmed this hypothesis [6]. Those right-handed neutrinos are known as sterile neutrinos [7, 8, 9, 10]. as follows [11, 12, 13]
$L_{iL}$ $=$ $\left(\begin{array}[]{c}\nu_{iL}\\ l_{iL}\end{array}\right)\sim\left({\bf 1},{\bf 2},-\frac{1}{2}\right),\,\ l_{iR}\sim\left({\bf 1},{\bf 1},-1\right),\,\ i=1,2,3.$ (4)
We show in parenthesis the transformations properties under the respective factors $(SU(3)_{C},SU(2)_{L},U(1)_{Y})$. The quark sector is defined as
$Q_{iL}$ $=$ $\left(\begin{array}[]{c}u_{iL}\\ d_{iL}\end{array}\right)\sim\left({\bf 3},{\bf 2},+\frac{1}{6}\right),\,\ u_{iR}\sim\left({\bf 3},{\bf 1},+\frac{2}{3}\right),\,\ d_{iR}\sim\left({\bf 3},{\bf 1},-\frac{1}{3}\right).$ (7)
In the SM, we have to introduce the following scalar field $\phi$ in the doublet representation of the $SU(2)_{L}$ [14, 15]
$\phi$ $=$ $\left(\begin{array}[]{c}\phi^{+}\\ \phi^{0}\end{array}\right)\sim\left({\bf 1},{\bf 2},+\frac{1}{2}\right),\,\ \Rightarrow\,\ \tilde{\phi}\equiv\left[\left(i\sigma_{2}\right)\phi^{*}\right]=\left(\begin{array}[]{c}\left(\phi^{0}\right)^{*}\\ -\phi^{-}\end{array}\right)\sim\left({\bf 1},{\bf 2},-\frac{1}{2}\right).$ (13)
This scalar fields get the following vacuum expectation value (VEV)
@@
The relevant terms in the SM lagrangian are given by
${\cal L}_{SM}={\cal L}^{Y}_{SM}-V(\phi)+{\cal L}^{kin}_{SM}.$ (20)
The charged fermions obtain their masses, via the following Yukawa coupling [11, 12, 13]
${\cal L}^{Y}_{SM}$ $=$ $\left[g^{l}_{ij}\left(\bar{L}_{iL}\phi\right)l_{jR}+g^{d}_{ij}\left(\bar{Q}_{iL}\phi\right)d_{jR}+g^{u}_{ij}\left(\bar{Q}_{iL}\tilde{\phi}\right)u_{jR}+hc\right].$ (21)
The scalar potential of the SM is defined in the following way [14, 15]
@@
Both the $\mu$-parameter and the $\lambda$-parameter are arbitrary and they are free parameters within the SM. For $\mu$-parameter, we can have the following two possibilities [14, 15]
  • $\mu^{2}>0:$ The gauge symmetry $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ is not broken;
  • $\mu^{2}<0:$ We break the $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ gauge symmetry to $SU(3)_{C}\times U(1)_{em}$.
When we impose the condition of extremes
$\left.\frac{\partial V(\phi)}{\partial\phi}\right|_{\langle\phi\rangle=v}=0,$ (23)
the minimum of the potential, in the case where $\mu^{2}<0$, is written in the following way
$v$$\equiv$ $\sqrt{\frac{-\mu^{2}}{\lambda}}=\frac{1}{\sqrt{\sqrt{2}G_{F}}}=246\,\ GeV,$
$G_{F}$ $=$ $\left(1.16639\pm 0.00002\right)\times 10^{-5}\,\ GeV^{-2}\approx\frac{1}{(293\,\ GeV)^{2}}.$ (24)
@@
$V(\phi^{\prime 0})$ $=$ $-\frac{\lambda}{4}v^{4},\,\ |\phi^{\prime 0}|\equiv\frac{v}{\sqrt{2}}=\sqrt{\frac{-\mu^{2}}{2\lambda}}>0.$ (26)
The field $h^{0}$ is CP even field and gets the following mass  [^2] A nice review about this interesting subject can be found in [15].
@@
We must emphasize, the SM do not predict any value for this mass, in fact it can get any values from 00 until a superior limit given by $M_{h^{0}}<1$ TeV [16, 17]. This mechanism is known as Brout-Englert-Higgs mechanism [15, 19].
The mass of the light Higgs boson $h^{0}$ was recently measured experimentally by ATLAS (A Toroidal LHC Apparatus) and CMS (Compact Muon Solenoid) collaborations, at Large Hadron Collider (LHC). Their measurements for the mass of the Higgs is [20, 21]
@@
After breaking the gauge symmetry the gauge boson get mass, through the following term
${\cal L}^{kin}_{SM}$ $=$ $\left(D_{m}\phi\right)^{\dagger}\left(D^{m}\phi\right),$
$D_{m}\phi$ $=$ $\left[\partial_{m}-ig\sum_{i=1}^{3}\left(\frac{\sigma^{i}}{2}\right)W^{i}_{m}-ig^{\prime}\left(\frac{1}{2}\right)W_{m}\right]\phi,$ (29)
where $W^{i}_{m}$ is the $SU(2)_{L}$ gauge bosons and $W_{m}$ is the $U(1)_{Y}$ gauge boson, while $g$ and $g^{\prime}$ are the gauge constant of the group $SU(2)_{L}$ and $U(1)_{Y}$, respectively. We obtain in addition to photon $A_{m}$, two massive bosons $W^{\pm}_{m}$ and $Z^{0}_{m}$ and their masses have the following expressions [^3] The numerical value of the first relation below was obtained from [22] [11, 12]
@@
$M^{2}_{Z^{0}}$ $=$ $\left(\frac{g^{2}+(g^{\prime})^{2}}{4}\right)v^{2}=\frac{g^{2}v^{2}}{4}\left(1+\tan^{2}\theta_{W}\right)=\frac{M^{2}_{W^{\pm}}}{\cos^{2}\theta_{W}},$
$W^{\pm}_{m}$ $=$ $\frac{1}{\sqrt{2}}\left(W^{1}_{m}\mp W^{1}_{m}\right),\,\ \left(\begin{array}[]{c}Z^{0}_{m}\\ A_{m}\end{array}\right)=\left(\begin{array}[]{cc}\cos\theta_{W}&-\sin\theta_{W}\\ \sin\theta_{W}&\cos\theta_{W}\end{array}\right)\left(\begin{array}[]{c}W_{3m}\\ W_{m}\end{array}\right),$ (36)
where $\theta_{W}$ is the Weinberg angle and it is defined as follows
@@
The average of the experimental values for the gauge bosons masses are [20, 22]
@@
Today we know neutrinos are massive particles [23, 24, 25, 26, 27, 28]. If we introduce sterile neutrinos in SM, $\nu_{iR}$, we can write the following Dirac mass term for neutrinos
${\cal L}^{\nu}_{\phi,L}=y^{\nu}_{ij}\left[\left(\bar{L}_{iL}\tilde{\phi}\right)\nu_{jR}+hc\right],$ (41)
where $\tilde{\phi}$ is defined in our Eq.(25). This mass term conserve lepton number defined in our Eq.(43) and in this case the neutrinos are Dirac-type particles. A priori, to solve this problem, we can also to introduce the following Majorana mass term for the left-handed neutrinos [25]
@@
where $B$ is the baryon number and $L_{e},L_{\mu}$ and $L_{\tau}$ are the lepton number of each lepton.
An interesting way to generate the Eq.(42), is to add the following term for the Yukawa coupling, defined by Eq.(21), lagrangian of SM [13]
$\overline{L^{c}_{iL}}L_{jL}\sim\left({\bf 1},{\bf 1\oplus 3},-1\right);\,\ L^{c}_{iL}=C\bar{\psi}^{T}\sim\gamma_{2}f^{*}=\left(\begin{array}[]{c}\nu^{c}_{iL}\\ l^{c}_{iL}\end{array}\right)\sim\left({\bf 1},{\bf 2},+\frac{1}{2}\right).$ (44)
for $f=\nu$ and $e$. Therefore, to provide mass to neutrinos, we need to introduce new fields within the SM context. We can add to the SM singlets and also the triplets, of scalars and/or fermions, in representation of $SU(2)_{L}$. When we introduce
@@
We can break the lepton number symmetry [^4] The total lepton number $L$ is defined in terms of $L_{e},L_{\mu}$ and $L_{\tau}$, see Eq.(43), in the following way $L=L_{e}+L_{\mu}+L_{\tau}$. when we introduce the new scalar $\Delta$ in the triplet representation [^5] The triplet $\Delta$ carries the lepton number of $2$., see Eq.(44), of the group $SU(2)_{L}$
$\Delta=\left(\begin{array}[]{cc}\frac{h^{+}}{\sqrt{2}}&H^{++}\\ \Delta^{0}&-\frac{h^{+}}{\sqrt{2}}\end{array}\right)\sim\left({\bf 1},{\bf 3},+1\right),\,\ \langle\Delta\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&0\\ V_{\Delta}&0\end{array}\right).$ (45)
@@
${\cal L}^{\nu}_{\Delta,L}$ $=$ $g^{\nu}_{ij}\left[\left(\overline{L^{c}_{iL}}\Delta L_{jL}\right)+hc\right]$
$=$ $g^{\nu}_{ij}\left[\overline{\nu^{c}_{iL}}\nu_{jL}\Delta^{0}+\left(\overline{\nu^{c}_{iL}}l_{jL}-\overline{l^{c}_{iL}}\nu_{jL}\right)\frac{h^{+}}{\sqrt{2}}+\overline{l^{c}_{iL}}l_{jL}H^{++}+hc\right],$
the Yukawa coupling $h_{ij}$ is the $3\times 3$ complex symmetric matrix. Those scalars $h^{+},\,\ H^{++}$ and $\Delta^{0}$ are known as bileptons [44]. The first term in the above equation is exactly our Eq.(42), where we can identify
@@
It is know as Gelmini-Roncadelli scheme [43], it is type II Seesaw mechanism [32, 33, 34, 35, 36] and neutrinos are Majorana-type particles rather than being a Dirac-type particle such as all charged fermions. The double beta decay experiments, is defined in the following way
@@
can show if the neutrinos are Majorana particles [45] or Dirac particles [46], see our Eq.(41), more information on this subject can be seen at [9, 10, 25, 47, 48, 49].
@@
Using the new scalar triplet, $\Delta$ defined in our Eq.(45), we can in addition to generating masses for neutrinos, jointly explain the recent measurement for $W$-boson mass, obtained by CDF collaboration [50]
$\left(M_{W^{\pm}}\right)_{CDF}=\left(80.4335\pm 0.0094\right)\,\ GeV,$ (51)
because the masses of the gauge bosons in the scheme of Gelmini-Roncadelli are written as follows [51, 52, 53]
$M^{2}_{W^{\pm}}$ $=$ $\frac{g^{2}}{4}\left(v^{2}+2V^{2}_{\Delta}\right)=\frac{g^{2}v^{2}}{4}\left(1+2R\right),\,\ R=\left(\frac{V_{\Delta}}{v}\right)^{2},$
@@
We can explain the new measurement for the $W$-boson mass if $V_{\Delta}\approx 10$ GeV, as can be easily seen in our Fig.(1).
Figure 1: We show the prediction about the behaviour for the gauge boson mass, $M_{W^{\pm}}\equiv M_{W}$, in the scheme of Gelmini-Roncadelli as a function of the Triplet VEV $V_{\Delta}$ (black curve) and we take $v=246$ GeV. The straight lines in red, green and blue are the values show in our Eqs.(LABEL:wsmval,39,51), respectively.
@@
Now, if we make the shift defined in our Eq.(25) together with the following expansion
@@
After carrying out the shift, defined in our Eqs.(25,55), we obtain the following mass matrices for the neutral scalars in the scheme of Gelmini-Roncadelli
@@
Where we get the following analytical results $det\left(M^{2}_{R}\right)\neq 0$ and $Tr\left(M^{2}_{R}\right)\neq 0$, therefore we have two massive scalars. The lightest Higgs, we will represent as $h^{0}$, and we also have another heavier Higgs, which we will denote as $H^{0}$. The mass of the lightest scalar, is the same as the Higgs boson of SM, compare with Eq.(27), is [^6] The complete expression for the mass of these scalars can be found in [57]. The mass of heavier Higgs, $H^{0}$, we will present below.
@@
The mass $M_{h^{0}}$ as function of the $\lambda_{1}$-parameter is shown in our Fig.(3). Where, we can see that we can get the central data of the lightest Higgs mass if we assume $\lambda_{1}\approx 0.1295$.
@@
We get two Goldstone bosons, because $det\left(M^{2}_{I}\right)=0$ and $Tr\left(M^{2}_{I}\right)=0$, one of them is the same $G^{0}$ obtained in SM. The extra Goldstone boson is the Majoron, $M^{0}\equiv I_{\Delta}$ [43]. The Majoron, $M^{0}$, can be detected experimentally through the following process [47, 48, 49]
@@
The existence of a non-massive Majoron, $M^{0}$, is excluded by experimental data [^7] Decay $Z^{0}\rightarrow M^{0}+X$. [58]. One possible solution for this problem, is to make the Majoron, $M^{0}$, invisible to the leptons and quarks, but we will not use this solution in this article.
There are another three possible solutions to remove the non-massive Majoron [56]
@@
Where $V_{G}(\phi,\Delta)$ is defined in our Eq.(54) and $\tilde{\phi}$ is defined in our Eq.(25). $\langle\Delta\rangle$ give mass only for neutrinos, review our Eq.(47), then it has tiny VEV, which is precisely the third case mentioned above. In the scheme of Ma, our Eq.(69) become
@@
The mass spectrum, we get beyond the lightest Higgs, $h^{0}$, given in Eq.(65), we obtain the heavier Higgs field, $H^{0}$, and its mass is given by
@@
Figure 4: We show the prediction about the behaviour for the heavier Higgs, $M_{H^{0}}$, in the scheme of Ma, see our Eq.(70), as a function of the Triplet VEV $V_{\Delta}$, and $v=246$ GeV. The values of $\mu\equiv\mu_{M}$ show in the box in several colors. $M^{exp}_{H^{0}}$ is the lower experimental values given at our Eq.(71) in orange line.
@@
therefore $M_{A^{0}}>M_{H^{0}}$. We show $M_{A^{0}}$ as function of, the VEV, $V_{\Delta}$ in our Fig.(5) and again our results for the mass of $M_{A^{0}}$ is in agreement experimental limits for the pseudoscalar, $A^{0}$, given by the following constraint [59]
@@
if $V_{\Delta}\leq 3$ GeV and $\mu\geq 10^{0}$ GeV.
Figure 5: We show the prediction about the behaviour for the pseudoscalar, $M_{A^{0}}$, in the scheme of Ma, see our Eq.(72), as a function of the Triplet VEV $V_{\Delta}$, and $v=246$ GeV. The values of $\mu\equiv\mu_{M}$ show in the box in several colors. $M^{exp}_{A^{0}}$ is the lower experimental values given at our Eq.(73) in orange line.
@@
Our numerical results satisfy the following experimental constraints $M_{h^{\pm}}>155$ GeV [20] and $M_{h^{\pm\pm}}>445$ GeV [62]. We want to emphasize that the masses of the charged scalars for the case $\lambda_{5}<0$ are heavier than in the opposite case.
Figure 6: We show the prediction about the behaviour for the masses, $M_{H^{0}}$ (blue curve), and $v=246$ GeV. The values of $\mu\equiv\mu_{M}$ show in the box in several colors. The $M_{h^{\pm}}$ (green curve) and $M_{h^{\pm\pm}}$ (red curve) when $\lambda_{5}>0$.
Figure 7: We show the prediction about the behaviour for the masses $M_{h^{\pm}}$ (green curve) and $M_{h^{\pm\pm}}$ (red curve) when $\lambda_{5}<0$ and $M_{H^{0}}=654.150$ GeV. The values contained in this graph are in agreement with the analytical results obtained by Eqs.(75,LABEL:numvall5n).
Supersymmetry, it is more known as SUSY, arose in theoretical papers more than 30 years ago independently by Golfand and Likhtman [63], Volkov and Akulov [64] and Wess and Zumino [65]. The supersymmetry algebra was introduced in [63]  [^9] They also constructed the first four-dimensional field theory with supersymmetry, (massive) quantum electrodynamics (SQED) of spinors and scalars. We will call it as Golfand-Likhtman model as suggested by Shiffman in his article [66].. The Volkov-Akulov article, however, deals only with fermions and its name is “Is the Neutrino a Goldstone Particle? [^10] It is the base of supergravity.. For interested people interested in the history of supersymmetry see [67, 68] and for the earliest Supersymmetric Standard Model see [69, 70], while anyone interested, how to work with supersymmetry see the great review article [71].